# Weight Space Representation Learning with Neural Fields

### v1

@@ 8.5 Pipeline
For the diffusion model architecture, we use a standard Diffusion Transformer (DiT) peebles2022scalable for standalone MLP weights. For LoRA weights, we employ the hierarchical LoRA layer encoder described in Section 3.5 of the main paper. During training, we use the DDPM objective with a linear noise schedule from $\beta_{1}=10^{-4}$ to $\beta_{T}=2\times 10^{-2}$ over $T=500$ timesteps. For generation, we use DDIM sampling with 100 steps to efficiently sample new weight representations. The diffusion transformer has 2880 hidden size (i.e., the size of each token after linear projection or layer encoder), 12 layers, and 16 self-attention heads. We train the diffusion transformer with batch size of 256 and learning rate of $2\times 10^{-4}$ for $6000$ epochs until convergence.
@@ 9.1 Distributional Difference
These metrics operate on features extracted by deep learned models, as deep learned feature extractors project the data into semantically rich embedding spaces where distances better correlate with human perception of similarity. We employ these metrics in their mathematical form rather than using modality-specific implementations like FID heusel2017gans or KID binkowski2018demystifying, as this allows for consistent evaluation across different data modalities. For 2D images, we use CLIP radford2021learning as the feature extractor; for 3D shapes, we use a PointNet++ qi2017pointnet++.
@@
Fréchet Distance (FD) frechet1957distance measures the distance between two multivariate Gaussian distributions fitted to feature representations of generated and reference samples. Given feature representations from a pretrained network, we compute the mean $\boldsymbol{\mu}_{P}$ and covariance $\boldsymbol{\Sigma}_{P}$ for generated distribution $P$ and mean $\boldsymbol{\mu}_{Q}$ and covariance $\boldsymbol{\Sigma}_{Q}$ for reference distribution $Q$. The Fréchet Distance is then computed as:
$\text{FD}(P,Q)=\frac{1}{N_{\text{feature}}}[\|\boldsymbol{\mu}_{P}-\boldsymbol{\mu}_{Q}\|_{2}^{2}+\text{Tr}\left(\boldsymbol{\Sigma}_{P}+\boldsymbol{\Sigma}_{Q}-2(\boldsymbol{\Sigma}_{P}\boldsymbol{\Sigma}_{Q})^{1/2}\right)]\;.$
@@
$\text{MMD}(P,Q)=\mathbb{E}_{\mathbf{x},\mathbf{x}^{\prime}}[\psi(\mathbf{x},\mathbf{x}^{\prime})]$ $+\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}}[\psi(\mathbf{y},\mathbf{y}^{\prime})]-2\mathbb{E}_{\mathbf{x},\mathbf{y}}[\psi(\mathbf{x},\mathbf{y})],$
where $\mathbf{x},\mathbf{x}^{\prime}\sim P$ are samples from the generated distribution and $\mathbf{y},\mathbf{y}^{\prime}\sim Q$ are samples from the reference distribution. This metric does not make the multivariate Gaussian assumption and is reported to be more reliable and sample efficient jayasumana2024rethinking; binkowski2018demystifying. We calculate this metric with 2 types of kernels.
@@
    $\psi_{p}(\mathbf{x},\mathbf{y})=(\gamma_{p}\cdot\mathbf{x}^{T}\mathbf{y}+c)^{d}$ with degree $d=3$ and offset $c=1$. Following the practice of KID binkowski2018demystifying, we choose $\gamma_{p}=1/N_{\text{feature}}$.
@@ 9.2 Distance-based Metrics for 3D Shapes
For 3D shapes, we calculate distance-based metrics following erkocc2023hyperdiffusion; vahdat2022lion; luo2021diffusion; zhou20213d. We denote the distance function as $D(\mathbf{x},\mathbf{y})$ for the Chamfer Distance between two shapes $\mathbf{x}$ and $\mathbf{y}$. The metrics are defined as:
@@ 10 Ablation Study
11 Additional Qualitative Results on Generation

### v2

@@ 8.5 Pipeline
For the diffusion model architecture, we use a standard Diffusion Transformer (DiT) [31] for standalone MLP weights. For LoRA weights, we employ the hierarchical LoRA layer encoder described in Section 3.5 of the main paper. During training, we use the DDPM objective with a linear noise schedule from $\beta_{1}=10^{-4}$ to $\beta_{T}=2\times 10^{-2}$ over $T=500$ timesteps. For generation, we use DDIM sampling with 100 steps to efficiently sample new weight representations. The diffusion transformer has 2880 hidden size (i.e., the size of each token after linear projection or layer encoder), 12 layers, and 16 self-attention heads. We train the diffusion transformer with batch size of 256 and learning rate of $2\times 10^{-4}$ for $6000$ epochs until convergence.
@@ 9.1 Distributional Difference
These metrics operate on features extracted by deep learned models, as deep learned feature extractors project the data into semantically rich embedding spaces where distances better correlate with human perception of similarity. We employ these metrics in their mathematical form rather than using modality-specific implementations like FID [16] or KID [2], as this allows for consistent evaluation across different data modalities. For 2D images, we use CLIP [33] as the feature extractor; for 3D shapes, we use a PointNet++ [32].
@@
Fréchet Distance (FD) [12] measures the distance between two multivariate Gaussian distributions fitted to feature representations of generated and reference samples. Given feature representations from a pretrained network, we compute the mean $\boldsymbol{\mu}_{P}$ and covariance $\boldsymbol{\Sigma}_{P}$ for generated distribution $P$ and mean $\boldsymbol{\mu}_{Q}$ and covariance $\boldsymbol{\Sigma}_{Q}$ for reference distribution $Q$. The Fréchet Distance is then computed as:
$\text{FD}(P,Q)=\frac{1}{N_{\text{feature}}}[$ $\|\boldsymbol{\mu}_{P}-\boldsymbol{\mu}_{Q}\|_{2}^{2}+$
$\text{Tr}\left(\boldsymbol{\Sigma}_{P}+\boldsymbol{\Sigma}_{Q}-2(\boldsymbol{\Sigma}_{P}\boldsymbol{\Sigma}_{Q})^{1/2}\right)]\;.$
@@
$\text{MMD}(P,Q)=\mathbb{E}_{\mathbf{x},\mathbf{x}^{\prime}}[\psi(\mathbf{x},\mathbf{x}^{\prime})]$ $+\mathbb{E}_{\mathbf{y},\mathbf{y}^{\prime}}[\psi(\mathbf{y},\mathbf{y}^{\prime})]$
$-2\mathbb{E}_{\mathbf{x},\mathbf{y}}[\psi(\mathbf{x},\mathbf{y})],$
where $\mathbf{x},\mathbf{x}^{\prime}\sim P$ are samples from the generated distribution and $\mathbf{y},\mathbf{y}^{\prime}\sim Q$ are samples from the reference distribution. This metric does not make the multivariate Gaussian assumption and is reported to be more reliable and sample efficient [19, 2]. We calculate this metric with 2 types of kernels.
@@
    $\psi_{p}(\mathbf{x},\mathbf{y})=(\gamma_{p}\cdot\mathbf{x}^{T}\mathbf{y}+c)^{d}$ with degree $d=3$ and offset $c=1$. Following the practice of KID [2], we choose $\gamma_{p}=1/N_{\text{feature}}$.
@@ 9.2 Distance-based Metrics for 3D Shapes
For 3D shapes, we calculate distance-based metrics following [9, 37, 26, 47]. We denote the distance function as $D(\mathbf{x},\mathbf{y})$ for the Chamfer Distance between two shapes $\mathbf{x}$ and $\mathbf{y}$. The metrics are defined as:
@@ 10 Ablation Study
11 Weight Space Interpolation
Figure 8 visualizes linear interpolations between pairs of mLoRA weight representations on FFHQ. We linearly interpolate between two instances’ LoRA weight pairs $(\mathbf{A}_{1},\mathbf{B}_{1})$ and $(\mathbf{A}_{2},\mathbf{B}_{2})$, evaluating the resulting neural field at each interpolation step. While the interpolated weights do not always produce smooth, perceptually gradual transitions between instances (as would be expected from a continuous learned latent space), this does not undermine our claims about the quality of the weight space representations. Smooth interpolation is characteristic of continuous latent spaces specifically optimized for this property, such as VAE latent codes, but structured representations like VQ-VAE quantized codes and point-cloud latents also lack this property yet achieve strong generative performance [37]. Our experiments demonstrate that mLoRA weights support high-quality generation (Tables 2–3 in the main paper) and exhibit clear semantic structure for classification (Table 4, Figure 5 in the main paper), which are the primary criteria for effective weight space representations.
Figure 8: Weight space interpolation. Linear interpolation between pairs of mLoRA-Asym weight representations on FFHQ. Columns show the two endpoint instances (leftmost, rightmost) and intermediate interpolated reconstructions.
12 Additional Qualitative Results
@@
To verify that generated samples are novel rather than memorized reproductions of training data, we perform a CLIP-based nearest-neighbor analysis on FFHQ generations from the mLoRA-Asym configuration. For each generated image, we retrieve its nearest neighbor from the training set using CLIP feature similarity. Figure 9 shows paired comparisons (left: generated, right: nearest training neighbor). The generated images are visually distinct from their nearest training neighbors, confirming that the diffusion model generalizes rather than memorizes.
Figure 9: Novelty check. For each generated FFHQ image (left), we show its nearest neighbor from the training set retrieved by CLIP feature similarity (right). Generated samples are visually distinct from training data, confirming generalization rather than memorization.
@@
13 Limitation and Future Work
First, our approach requires all instances to share the same pre-trained base model and initialization. This is a practical limitation: the base model may not suit all INR architectures, and meaningful weight space comparisons between instances trained on different base models are not possible. Future work could explore methods to reduce this requirement or to align weight spaces across different base models.
Second, our approach requires finetuning a base-model which is computationally more expensive than fitting small MLPs. This requirement prevents evaluation on datasets with hundreds of thousands of samples, constraining our experiments to datasets with thousands of instances. Future work could explore methods to eliminate this requirement, or develop more computationally efficient multiplicative LoRA adaptation procedures.